Hi Can I have some help with a linear function is given?
g(x) = -7x + 8
(a) Find the average rate of change of the function between x = a and x = a + h.

(b) Find the slope of the line

Answer by stanbon(75887)

g(x) = -7x + 8
(a) Find the average rate of change of the function between x = a
and x = a + h.
-----
Average rate = [g(a+h)-g(a)/[(a+h)-a]= [g(a+h)-g(a)]/h
-------------
g(a+h) = -7(a+h)+8 = -7a-7h+8
g(a) = -7a+8
----------
Average rate =[(-7a-7h+8)-(-7a+8)]/h = -7h/h = -7
----------------------------

(b) Find the slope of the line
Since g(x) =-7x+8, the slope is -7
==================
Cheers,
Stan H.